A rectangular bedroom is 2 ft longer than it is wide. Its area is 288 ft2. What is the width of the room?
Answer by algebrahouse.com(1659)

x = width
x + 2 = length {the room is 2 ft longer than it is wide}

Area of a rectangle = width x length

x(x + 2) = 288 {substituted dimensions into area formula}
x² + 2x = 288 {used distributive property}
x² + 2x - 288 = 0 {subtracted 288 from each side}
(x + 18)(x - 16) = 0 {factored into two binomials}
x + 18 = 0 or x - 16 = 0 {set each factor equal to 0}
x = -18 or x = 16 {solved each equation for x}
x = 16 {width cannot be negative}

width is 16 ft

For more help from me, visit: www.algebrahouse.com